what is the present value of
6,000$ in 10 years at 6%
26,000$ in 5 years at 12%
15,000$ in 15 years at 8%?
Answer #1

Present Value=Future Value/(1+rate)^t

1.
=6000/1.06^10=3350.368661

2.
=26000/1.12^5=14753.09825

3.
=15000/1.08^15=4728.625574